# ARR: Question Answering with Large Language Models via Analyzing, Retrieving, and Reasoning

### v2

@@ 6 Generalizability
Size Method BBH MMLU MMLU-Pro Avg.
1B Baseline 35.88 43.27 21.62 33.59
CoT 36.30 41.10 22.74 33.38
ARR 39.02 42.70 23.49 35.07
3B Baseline 45.65 48.26 30.88 41.60
CoT 46.89 46.80 30.03 41.24
ARR 51.97 52.82 33.39 46.06
8B Baseline 52.19 60.68 38.75 50.54
@@
Table 6: Model size experiments. The zero-shot performance (Accuracy %) of LLaMA3-Chat models of different sizes on multiple-choice QA datasets.
Figure 3: Model size experiments. The trend of QA performance changes as the model becomes larger.
6.1 Model Sizes
We evaluate the LLaMA3-Chat models of different sizes, i.e., 1B, 3B, and 8B (default) parameters, on multiple-choice QA tasks. As the accuracy scores (%) shown in Table 6, our ARR method brings solid performance gains over the Baseline method and consistently outperforms zero-shot CoT. For the 1B model, ARR slightly underperforms the Baseline method on MMLU, likely due to the weaker instruction-following ability in smaller models. Still, our ARR method achieves overall performance improvements over the Baseline in the 1B setting. In addition, Figure 3 illustrates the trend of QA performance changes as the model becomes larger. The results conform to the scaling laws of language models Kaplan et al. (2020), demonstrating the potential of the proposed ARR method when applied to larger models.
Series Method BBH MMLU MMLU-Pro Avg.
Qwen Baseline 39.21 48.36 32.35 39.97
CoT 36.66 44.91 29.26 36.94
ARR 40.50 50.34 39.10 43.31
Gemma Baseline 40.09 45.46 23.45 36.33
CoT 44.39 47.17 26.20 39.25
ARR 45.31 50.73 26.98 41.01
Mistral Baseline 46.27 55.61 30.68 44.19
CoT 53.42 61.16 34.73 49.77
ARR 53.55 61.49 35.21 50.08
Table 7: LLM series experiments. The zero-shot performance (Accuracy %) of 7B-Chat models of different LLM series on multiple-choice QA datasets.
6.2 LLM Series
To verify the effectiveness of our ARR method on models other than LLaMA3 Dubey et al. (2024), we conduct experiments on 7B-Chat LLMs of different series: Qwen2.5 Yang et al. (2024), Gemma Team et al. (2024b, c), and Mistral Jiang et al. (2023). The results in Table 7 exhibit a consistent superiority of the proposed ARR method over the Baseline and CoT methods. This is similar to the findings in the main experiments (Table 2), solidifying the efficacy and generalizability of ARR.
Temp. Method BBH MMLU MMLU-Pro Avg.
0.0 Baseline 52.19 60.68 38.75 50.54
@@ 6.2 LLM Series
0.5 Baseline 50.19 59.35 36.88 48.81
CoT 56.58 60.82 37.82 51.74
ARR 58.87 62.87 42.64 54.79
1.0 Baseline 46.33 54.80 33.10 44.74
CoT 51.46 55.57 33.00 46.68
ARR 52.90 56.58 36.73 48.74
1.5 Baseline 40.84 45.03 26.85 37.57
CoT 42.53 44.85 25.61 37.66
ARR 42.65 45.16 27.44 38.42
Table 8: Generation temperature experiments. The zero-shot performance (Accuracy %) of the LLaMA3-8B-Chat model on multiple-choice QA datasets using different generation temperatures (default: 0.0).
6.3 Generation Temperatures
As shown in Table 8, our ARR method surpasses the Baseline and CoT methods with different temperatures, demonstrating a strong robustness of ARR. In addition, we observe that the model generally performs better when the temperature is lower.
6.4 Few-shot Generation
Few-shot Examples with Rationales.
For each subtask in a QA dataset, we randomly pick $10$ examples from the training or validation set if they exists. If a subtask only has the test set, $10$ test examples are held out for few-shot usage, slightly reducing the number of items for evaluation. For each raw example, we construct the CoT and ARR rationales using GPT-4o Hurst et al. (2024). Specifically, the input prompts provided to GPT-4o match those used in the evaluation experiments under CoT/ARR settings. The model’s output is extracted as CoT/ARR rationales. In few-shot examples, these rationales, along with correct answers, are appended to the answer trigger sentence $\phi$. For the Baseline setting, few-shot examples include correct answers for in-context learning (ICL) Brown et al. (2020); Dong et al. (2024) but exclude rationales.
Shot Method BBH MMLU MMLU-Pro Avg.
0 Baseline 52.19 60.68 38.75 50.54
CoT 58.40 62.08 40.10 53.53
ARR 59.01 63.51 42.72 55.08
1 Baseline 35.68 44.80 28.62 36.37
@@ 6.4 Few-shot Generation
Table 9: Few-shot experiments. The few-shot performance (Accuracy %) of the LLaMA3-8B-Chat model on multiple-choice QA datasets using 1, 3, and 5 few-show examples with rationales.
@@
Table 9 presents the accuracy scores (%) of the LLaMA3-Chat model on multiple-choice QA tasks. Using different numbers of few-shot examples (1, 3, and 5), our few-shot ARR method outperforms the ICL Baseline and few-shot CoT Wei et al. (2022) methods on average.
@@ 7 Conclusion
In this work, we introduce ARR, an intuitive, simple, and general prompting method that effectively enhances the question-answering performance of LLMs by integrating three key steps: analyzing the question’s intent, retrieving relevant information, and reasoning step by step. Extensive experiments across diverse QA tasks demonstrate that ARR consistently improves the Baseline (without ARR prompting) and outperforms the CoT prompting method. Ablation and case studies further validate the positive contributions of each component, with intent analysis proving particularly crucial. In addition, evaluations across various model sizes, LLM series, and generation configurations confirm the effectiveness, robustness, and generalizability of the proposed ARR method.
@@
We did not explore variations or paraphrases of the proposed ARR prompts, opting instead for the straightforward expressions presented in this paper. While certain phrasings may further enhance performance, the core idea remains the same.
In addition, resource constraints limited our focus to open-weights LLMs with no more than 8B parameters. However, the results from model size experiments (§ 6.1) align with the scaling laws for language models Kaplan et al. (2020), demonstrating the potential and generalizability of our ARR method when applied to larger models.
Lastly, we observe that some generated rationales in the Reasoning Generation stage are repetitive and redundant. A dynamic stopping strategy or post-processing to filter out redundancies can reduce the computational cost and potentially further boost the final QA accuracy.
@@ A.4 Reproducibility
For the reproducibility of this work, we set the generation temperature as 00 by default and disable token sampling for deterministic generation. In addition, we pre-set the random seed for all random modules at the beginning of each experiment session. By an unofficial tradition666 “The answer to the ultimate question of life, the universe, and everything is forty-two.” Adams (1979), we set $42$ as the random seed and do not tune the value. To validate the reproducibility, we ran the main experiments twice and obtained the same results as shown in Table 2. Our source code is available on GitHub: https://github.com/YuweiYin/ARR
@@ A.5 Experimental Cost
In the reasoning generation stage, the total computational cost is approximately 6,900 GPU hours on NVIDIA V100 clusters (about 288 days) and 1,300 hours on A100 clusters (about 54 days). We only use V100 clusters for option selection, and the overall running time is approximately 500 hours (about 21 days). In the few-shot experiments, the expense for GPT-4o API calls to generate CoT and ARR rationales for few-shot examples is below US\$20.
Appendix B More Cases
As discussed in § 5.3, this section presents additional cases (Table 1221) from our main experiments (§ 5, Table 2) to provide a closer look at how the ARR method enhances reasoning and improves QA accuracy. We present one case for each QA dataset introduced in § 4.1.
@@ Appendix B More Cases
Table 12: Case study. An instance from the Boolq dataset, where our ARR method correctly answers the question but the Baseline and CoT methods fail.
@@
Table 13: Case study. An instance from the LogiQA dataset, where our ARR method correctly answers the question but the Baseline and CoT methods fail. Analysis: the correct station order can be Jiushang-Yinling-Haoyun-Fuyi-Huyao or Yinling-Jiushang-Haoyun-Fuyi-Huyao.
@@
Table 14: Case study. An instance from the Commonsense QA dataset, where our ARR method correctly answers the question but the Baseline and CoT methods fail.
@@
Table 15: Case study. An instance from the SocialIQA dataset, where our ARR method correctly answers the question but the Baseline and CoT methods fail. Analysis: Option B is not wrong, but Option A is more appropriate as they are in a romantic relationship (“significant other”).
@@
Table 16: Case study. An instance from the SciQ dataset, where our ARR method correctly answers the question but the Baseline and CoT methods fail.
@@
Table 17: Case study. An instance from the OpenBookQA dataset, where our ARR method correctly answers the question but the Baseline and CoT methods fail. Analysis: The CoT method provides a correct fact about the direct relationship (“A direct relationship means that $\dots$”), but wrongly infers that “this is a direct relationship” from “an increase in friction leads to a decrease in speed.” This is because the wrong statement is made without conditioning on the correct fact that is presented after the statement (due to the autoregressive generation). In contrast, the ARR method performs reasoning based on sufficient context after analyzing the intent and finding relevant information.
@@
Table 18: Case study. An instance from the ARC dataset, where our ARR method correctly answers the question but the Baseline and CoT methods fail.
@@
Table 19: Case study. An instance from the BBH dataset, where our ARR method correctly answers the question but the Baseline and CoT methods fail. Analysis: The Baseline method rushes to conclusions and then tries to present an explanation. However, the false statement confuses the whole reasoning process and makes the model caught in an infinite self-doubting loop.
@@
Table 20: Case study. An instance from the MMLU dataset, where our ARR method correctly answers the question but the Baseline and CoT methods fail. Analysis: The Baseline and CoT methods jump to the wrong conclusions confidently, while the ARR method understands the question first and analyzes each option correctly.
@@
Table 21: Case study. An instance from the MMLU-Pro dataset, where our ARR method correctly answers the question but the Baseline and CoT methods fail. Analysis: The ARR method grasps the question’s intent and implications before answering.

### v3

@@ 6 Generalizability
Size Method BBH MMLU MMLU-Pro Avg. $\Delta$
1B DA 35.88 43.27 21.62 33.59 0
CoT 36.30 41.10 22.74 33.38 --0.21
ARR 39.02 42.70 23.49 35.07 $+$1.48
3B DA 45.65 48.26 30.88 41.60 0
CoT 46.89 46.80 30.03 41.24 --0.36
ARR 51.97 52.82 33.39 46.06 $+$4.46
8B DA 52.19 60.68 38.75 50.54 0
CoT 58.40 62.08 40.10 53.53 $+$2.99
ARR 59.01 63.51 42.72 55.08 $+$4.54
Table 5: Model size experiments. The zero-shot performance (Accuracy %) of LLaMA3-Chat models of different sizes on multitask QA datasets.
6.1 Model Sizes
We evaluate the LLaMA3-Chat models of different sizes, i.e., 1B, 3B, and 8B (default) parameters, on multitask QA tasks. As the accuracy scores (%) shown in Table 5, our ARR method brings solid performance gains over the DA method and consistently outperforms zero-shot CoT. For the 1B model, ARR slightly underperforms the DA method on MMLU, likely due to the weaker instruction-following ability in smaller models. Still, our ARR method achieves overall performance improvements over the DA in the 1B setting. Observing the improvements over the DA method, larger models benefit more from ARR. The results conform to the scaling laws of language models Kaplan et al. (2020), demonstrating the potential of ARR when applied to larger models.
Series Method BBH MMLU MMLU-Pro Avg.
Qwen DA 39.21 48.36 32.35 39.97
CoT 36.66 44.91 29.26 36.94
ARR 40.50 50.34 39.10 43.31
Gemma DA 40.09 45.46 23.45 36.33
CoT 44.39 47.17 26.20 39.25
ARR 45.31 50.73 26.98 41.01
Mistral DA 46.27 55.61 30.68 44.19
CoT 53.42 61.16 34.73 49.77
ARR 53.55 61.49 35.21 50.08
Table 6: LLM series experiments. The zero-shot performance (Accuracy %) of 7B-Chat models of different LLM series on multitask QA datasets.
6.2 LLM Series
To verify the effectiveness of our ARR method on open models other than LLaMA3 Dubey et al. (2024), we conduct experiments on 7B-Chat LLMs of different series: Qwen2.5 Yang et al. (2024), Gemma Team et al. (2024b, c), and Mistral Jiang et al. (2023). The results in Table 6 exhibit a consistent superiority of the proposed ARR method over the baseline methods. This is similar to the findings in the main experiments (Table 2), solidifying the efficacy and generalizability of ARR.
Method OBQA MMLU-sub
DA 98.00 88.55
CoT 97.20 88.55
ARR 98.20 88.91
Table 7: Proprietary LLM experiments. The zero-shot performance (Accuracy %) of GPT-4o on the OpenBookQA (OBQA) dataset and a subset of MMLU.
6.3 Proprietary LLMs
To further validate the efficacy of our ARR method on proprietary LLMs Hurst et al. (2024); Anthropic (2024); Team et al. (2024a) with a humongous number of parameters and state-of-the-art (SOTA) performances, we conduct experiments using GPT-4o Hurst et al. (2024) on OBQA (involving World Knowledge) and MMLU (involving Multitask Understanding). As MMLU is a large benchmark with 50+ subtasks, we sample 10 instances from each subtask to form an “MMLU-sub” subset with 500+ items for GPT evaluation.
As presented in Table 7, our ARR framework continues to yield performance gains even when applied to SOTA models like GPT-4o. While the improvement is not as pronounced as for smaller LLMs—likely due to the already excellent base performance—ARR still provides measurable benefits. In contrast, the CoT method either diminishes GPT-4o’s performance or offers negligible gains. This suggests that GPT-4o may already engage in CoT-like reasoning intrinsically, yet lacks the structured three-step process introduced by ARR.
Temp. Method BBH MMLU MMLU-Pro Avg.
0.0 DA 52.19 60.68 38.75 50.54
@@ 6.3 Proprietary LLMs
0.5 DA 50.19 59.35 36.88 48.81
CoT 56.58 60.82 37.82 51.74
ARR 58.87 62.87 42.64 54.79
1.0 DA 46.33 54.80 33.10 44.74
CoT 51.46 55.57 33.00 46.68
ARR 52.90 56.58 36.73 48.74
1.5 DA 40.84 45.03 26.85 37.57
CoT 42.53 44.85 25.61 37.66
ARR 42.65 45.16 27.44 38.42
Table 8: Generation temperature experiments. The zero-shot performance (Accuracy %) of the LLaMA3-8B-Chat model on multitask QA datasets using different generation temperatures (default: 0.0).
6.4 Generation Temperatures
As shown in Table 8, our ARR method surpasses the baseline methods with different temperatures, demonstrating a strong robustness of ARR. In addition, we observe that the model generally performs better when the temperature is lower.
6.5 Few-shot Generation
Few-shot Examples with Rationales.
For each subtask in a QA dataset, we randomly pick $10$ examples from the training or validation set if they exists. If a subtask only has the test set, $10$ test examples are held out for few-shot usage, slightly reducing the number of items for evaluation. For each raw example, we construct the CoT and ARR rationales using GPT-4o Hurst et al. (2024). Specifically, the input prompts provided to GPT-4o match those used in the evaluation experiments under CoT/ARR settings. The model’s output is extracted as CoT/ARR rationales. In few-shot examples, these rationales, along with correct answers, are appended to the answer trigger sentence $\phi$. For the Direct Answer (DA) setting, few-shot examples include correct answers for in-context learning (ICL) Brown et al. (2020); Dong et al. (2024) but exclude rationales.
Shot Method BBH MMLU MMLU-Pro Avg.
0 DA 52.19 60.68 38.75 50.54
@@ 6.5 Few-shot Generation
1 DA 35.68 44.80 28.62 36.37
@@
Table 9: Few-shot experiments. The few-shot performance (Accuracy %) of the LLaMA3-8B-Chat model on multitask QA datasets using 1, 3, and 5 few-show examples with rationales.
@@
Table 9 presents the accuracy scores (%) of the LLaMA3-Chat model on multitask QA tasks. Using different numbers of few-shot examples (1, 3, and 5), our few-shot ARR method outperforms the DA (i.e., ICL) and few-shot CoT Wei et al. (2022) methods on average.
@@ 7 Conclusion
In this work, we introduce ARR, an intuitive, effective, and general QA framework that effectively enhances the question-answering performance of LLMs by integrating three key steps: analyzing the question’s intent, retrieving relevant information, and reasoning step by step. Extensive experiments across diverse QA tasks demonstrate that ARR consistently outperforms baseline methods including Direct Answer (DA) and Chain-of-Thought (CoT) prompting. Ablation and case studies further validate the positive contributions of each component, with intent analysis proving particularly crucial. Furthermore, experiments on ARR prompt variations indicate that ARR remains effective regardless of the specific prompt implementation. In addition, evaluations across various model sizes, LLM series, and generation configurations confirm the effectiveness, robustness, and generalizability of the proposed ARR method.
@@
Resource constraints limited our focus to open-weights LLMs with no more than 8B parameters. However, the results from model size experiments (§ 6.1) align with the scaling laws for language models Kaplan et al. (2020), demonstrating the potential and generalizability of our ARR method when applied to larger models. Moreover, § 6.3 demonstrate that ARR continues to yield performance gains even when applied to SOTA models like GPT-4o.
As a general framework, ARR can be implemented in different ways. In this work, we realize ARR as a test-time prompting method because this is the most natural approach to elicit the power of pre-trained foundation models like LLMs. Section 5.3 shows that ARR outperforms baselines regardless of the specific prompt design, showing the efficacy of this three-step solution. Beyond prompting, the effectiveness of ARR sheds light on other approaches that may enhance LLMs, e.g., incorporating the three-step recipe of ARR (especially intent analysis) for LLM post-training.
Lastly, as mentioned in § 3.1, we only consider $10$ multiple-choice question answering (MCQA) tasks, where the model is asked to answer the question by selecting an option from a list of choices. However, the ARR outputs of the Reasoning Generation stage (§ 3.2) can be used for QA tasks requiring free-form generation. As studied in recent research (Yin et al., 2025, Table 3), ARR works excellently on multiple mathematical reasoning benchmarks. Such tasks are not MCQA, and there is no “Option Selection” stage.
@@ A.4 Reproducibility
For the reproducibility of this work, we set the generation temperature as 00 by default and disable token sampling for deterministic generation. In addition, we pre-set the random seed for all random modules at the beginning of each experiment session. By an unofficial tradition777 “The answer to the ultimate question of life, the universe, and everything is forty-two.” Adams (1979), we set $42$ as the random seed and do not tune the value. To validate the reproducibility, we ran the main experiments twice and obtained the same results as shown in Table 2. Our source code is available on GitHub: https://github.com/YuweiYin/ARR
@@ A.5 Experimental Cost
In the reasoning generation stage, the total computational cost is approximately 8,000 GPU hours on NVIDIA V100 clusters (about 333 days) and 1,300 hours on A100 clusters (about 54 days). We only use V100 clusters for option selection, and the overall running time is approximately 560 hours (about 23 days). The expense for GPT-4o API calls in § 6.3 and § 6.5 is below US\$100.
BoolQ LogiQA CSQA SIQA SciQ OBQA ARC BBH MMLU MMLU-Pro
w/o RG 77.86 35.64 50.37 47.49 91.20 69.80 64.61 50.26 45.54 29.60 56.24
DA 84.16 35.79 72.97 69.55 85.90 72.20 82.59 52.19 60.68 38.75 65.48
CoT 84.65 38.10 73.71 68.12 93.70 78.20 84.31 58.40 62.08 40.10 68.14
ARR 86.33 39.02 74.94 70.98 94.40 80.00 84.84 59.01 63.51 42.72 69.58
V1 84.40 36.56 75.51 68.63 93.10 83.20 84.05 61.58 63.45 42.96 69.34
V2 85.14 37.63 76.82 70.68 93.90 82.80 85.90 60.66 65.53 44.91 70.40
V3 84.68 38.71 75.76 69.34 93.70 83.40 85.92 59.59 65.05 44.31 70.05
V4 84.40 39.94 77.31 68.78 93.90 84.00 87.01 63.08 65.42 44.38 70.82
V5 84.22 38.10 76.25 69.34 93.40 81.40 85.36 61.02 64.61 44.12 69.78
Vars AVG 84.57 38.19 76.33 69.35 93.60 82.96 85.65 61.19 64.81 44.14 70.08
Vars STD $\pm$0.32 $\pm$1.12 $\pm$0.66 $\pm$0.72 $\pm$0.31 $\pm$0.87 $\pm$0.96 $\pm$1.15 $\pm$0.75 $\pm$0.64 $\pm$0.75
Table 12: Main experiments and ARR prompt variation experiments. “ARR” is the original prompt design, and “V1”–“V5” are five paraphrased prompt variants. “Vars AVG” and “Vars STD” denote the average and standard divination of the accuracy score on each QA dataset, respectively.
Appendix B ARR Prompt Variants
As mentioned in § 5.3, we conduct experiments on different ARR prompt variants to show that ARR works effectively irrespective of specific prompt design. The original ARR prompt is paraphrased into five different versions by GPT-4o Hurst et al. (2024). The full experimental results on 10 QA datasets are presented in Table 12.
  • V1:
    Answer: Let’s identify the question’s intent, gather the necessary information, and then work through a logical, step-by-step solution.
  • V2:
    Answer: We’ll begin by examining what the question is asking, then retrieve any relevant details, and finally provide a well-reasoned answer step by step.
  • V3:
    Answer: First, we’ll interpret the purpose behind the question, collect supporting information, and proceed to solve it methodically.
  • V4:
    Answer: Let’s break this down by understanding the goal of the question, pulling in the required data, and then reasoning through the answer in a clear sequence.
  • V5:
    Answer: To approach this, we’ll clarify the question’s intent, locate pertinent information, and then build our answer using structured, logical reasoning.
Appendix C Case Study
As mentioned in § 5.1, this section presents case studies (Table 1322) from our main experiments (§ 5, Table 2) to provide insights into how the ARR method enhances reasoning and improves QA accuracy. We present one case for each QA dataset introduced in § 4.1.
@@ Appendix C Case Study
Table 13: Case study. An instance from the Boolq dataset, where our ARR method correctly answers the question but the baseline methods fail.
@@
Table 14: Case study. An instance from the LogiQA dataset, where our ARR method correctly answers the question but the baseline methods fail. Analysis: the correct station order can be Jiushang-Yinling-Haoyun-Fuyi-Huyao or Yinling-Jiushang-Haoyun-Fuyi-Huyao.
@@
Table 15: Case study. An instance from the Commonsense QA dataset, where our ARR method correctly answers the question but the baseline methods fail.
@@
Table 16: Case study. An instance from the SocialIQA dataset, where our ARR method correctly answers the question but the baseline methods fail. Analysis: Option B is not wrong, but Option A is more appropriate as they are in a romantic relationship (“significant other”).
@@
Table 17: Case study. An instance from the SciQ dataset, where our ARR method correctly answers the question but the baseline methods fail.
@@
Table 18: Case study. An instance from the OpenBookQA dataset, where our ARR method correctly answers the question but the baseline methods fail. Analysis: The CoT method provides a correct fact about the direct relationship (“A direct relationship means that $\dots$”), but wrongly infers that “this is a direct relationship” from “an increase in friction leads to a decrease in speed.” This is because the wrong statement is made without conditioning on the correct fact that is presented after the statement (due to the autoregressive generation). In contrast, the ARR method performs reasoning based on sufficient context after analyzing the intent and finding relevant information.
@@
Table 19: Case study. An instance from the ARC dataset, where our ARR method correctly answers the question but the baseline methods fail.
@@
Table 20: Case study. An instance from the BBH dataset, where our ARR method correctly answers the question but the baseline methods fail. Analysis: The DA (Direct Answer) method rushes to conclusions and then tries to present an explanation. However, the false statement confuses the whole reasoning process and makes the model caught in an infinite self-doubting loop.
@@
Table 21: Case study. An instance from the MMLU dataset, where our ARR method correctly answers the question but the baseline methods fail. Analysis: The DA (Direct Answer) and CoT methods jump to the wrong conclusions confidently, while the ARR method understands the question first and analyzes each option correctly.
@@
Table 22: Case study. An instance from the MMLU-Pro dataset, where our ARR method correctly answers the question but the baseline methods fail. Analysis: The ARR method grasps the question’s intent and implications before answering.